# An Information-Theoretic Framework for Receiver Quantization in Communication

### v1

@@ VII-B On Improving Uniform Quantization by Post-Processing
In [57, 39], it has been shown that, the GMI (14) given in Proposition 1 can be improved by introducing a post-processing of the channel output as $Y\to\tilde{Y}$, thereby extending the channel $X\to Y$ to $X\to\tilde{Y}$; The GMI is maximized by letting
@@
i.e., using a conditional expectation operator as the post-processor, which gives an MMSE estimation of $X$ under the observation $Y$. Thus we can infer that the optimized quantizer with equispaced thresholds considered in Sec. VII-A is equivalent to a uniform quantizer combined with an MMSE post-processing of its output. In fact, for a quantizer with $\ell$-spaced thresholds, the optimal levels have been shown to be [30]
$y^{*}_{k}=\frac{\phi((k-1)\ell)-\phi(k\ell)}{Q((k-1)\ell)-Q(k\ell)},\;k=1,...,K-1$ (151)
@@
$y^{*}_{K}=\frac{\phi((K-1)\ell)}{Q((K-1)\ell)}.$ (152)
Such levels satisfy $y^{*}_{k}=\mathrm{E}\left[V|V\in[(k-1)\ell,k\ell)\right],\;k=1,...,K-1$ and $y^{*}_{K}=\mathrm{E}\left[V|V\geq(K-1)\ell\right]$, where $V\sim\mathcal{N}(0,\sigma_{v}^{2})$ is the input of the quantizer; i.e., the optimal level $y^{*}_{k}$ is the centroid of the corresponding interval. This is consistent with the aforementioned result that a conditional expectation operator maximizes the GMI. Comparing Fig. 17(e) and Fig. 17(d), we see that, under the same loading factor $L=k\ell$, such a post-processing indeed improves the GMI slightly. Interestingly, the optimal loading factor that maximizes the GMI in Fig. 17(e) is smaller than the optimal loading factor in Fig. 17(d). In other words, when post-processing is available, the optimal step size of the uniform quantizer becomes smaller, and a concatenation of an MMSE uniform quantizer and an MMSE post-processor is strictly worse than the optimized quantizer with equispaced thresholds.
Since the example Fig. 17(e) shows that the gain of the MMSE post-processing is limited for low-resolution uniform quantization, we may infer that the gain is also limited under high resolution. In fact, as the step size decreases, the centroid of the interval $[(k-1)\ell,k\ell)$ given in (151) tends to be its midpoint. [^11] In [54] the asymptotic convergence of centroids to midpoints was proved under uniform quantization with infinitely many levels. To see this, note that for an arbitrary $c>\ell/2$,
@@
=\;= $\lim\limits_{\ell\to 0}\frac{\frac{c-\frac{\ell}{2}}{2}\phi\left(c-\frac{\ell}{2}\right)+\frac{c+\frac{\ell}{2}}{2}\phi\left(c+\frac{\ell}{2}\right)}{\frac{1}{2}\left(\phi\left(c+\frac{\ell}{2}\right)+\phi\left(c-\frac{\ell}{2}\right)\right)}$ (153b)
=\;= $c,$ (153c)
which follows from L’Hôpital’s rule. Moreover, according to (102), the largest level $y^{*}_{K}$ given in (152) satisfies
$(K-1)\ell>y_{K}^{*}>(K-1)\ell+\frac{1}{(K-1)\ell},$ (154)
implying that $y_{K}^{*}/\ell$ converges to $K-1$ as $\ell\to 0$. From these facts we can conclude that the gain of MMSE post-processing asymptotically vanishes as the resolution increases.
VII-C On Possible Gain of Futher Quantization Rules
@@
As shown in Fig. 17(b), for nonuniform quantizer, there is also a consistency between GMI maximization and MSE minimization, but we must optimize all the thresholds and levels, rather than only a single factor ($g$ or $L$). In [References, Appendix E] a proof of this consistency has been given. Thus, the maximum achievable rate can be expressed as
@@
where $\mathsf{mmse}^{\textrm{nu}}$ is the normalized minimum MSE of all possible nonuniform quantizers. This result can also be understood via a similar geometrical interpretation as that given in Sec. IV-C. From (155) we can infer that, under our setting of transceiver architecture, the maximum gain in achievable rate from replacing the optimal uniform quantizer by the optimal nonuniform quantizer is determined by the reduction of the normalized MMSE from $\mathsf{mmse}$ to $\mathsf{mmse}^{\textrm{nu}}$. The numerical results in Fig. 17(a) and Fig. 17(b) show that the rate gain is marginal for $b=2$. For higher resolutions the rate gain grows slowly. When $b=4$, according to the numerical results in [30], the optimal non-uniform quantization improves the saturation rate $\log(1/\gamma)$ by no more than $4.4\%$, and optimized quantization with equispaced thresholds improves the saturation rate by $1.5\%$. In fact, for nonuniform quantization with Gaussian input, the asymptotic quantization theory, in particular the Panter-Dite formula [73], yields a “6-dB-per-bit” rule [^12] There is another version of the $6$-dB-per-bit rule as $B_{\mathrm{eff}}=(\mathsf{SNR}_{q}-1.76)/6.02$, which has been widely used in practice to calculate the effective number of bits (ENoB) of a quantizer [2, 3]. It is derived under assumptions that the granular distortion is uniformly distributed over $[-\ell/2,\ell/2]$ and there is no overload distortion; see [2] for more details. for the maximum SNR as [41]
@@
where $\mathsf{SNR}_{q}^{\textrm{nu}}=1/\mathsf{mmse}^{\textrm{nu}}$. Therefore, under our setting, the rate gain increases logarithmically with the resolution $b$, and for $b\leq 10$ the gain is limited (cf. (141)).
According to (156), even if the best nonuniform scalar quantization is used, the MSE achieved in the high-resolution regime is still $2.72$ times larger than the theoretical limit implied by the rate-distortion function of a Gaussian source [42]. [^13] The comparison is made under the assumption that the signal at the receiver front-end is represented by $b$ bits per channel use before feeding into the decoder. The MSE of the representation can be reduced by lossy source coding (treating the received signal as a source). A scalar quantizer is a very simple lossy source coding scheme. If we allow coded uniform scalar quantization (i.e., representing the quantizer output by a variable-rate lossless code), the gap can be reduced to only $1.53$ dB ($\pi e/6$) worse than the theoretical limit (see [41] and references therein). Unfortunately, coded quantization does not help in communication receivers since the bottleneck therein is the limited resolution of the quantizer rather than the cost of representing its output. Applying vector quantization to communication receivers is a more challenging topic. Nevertheless, the gap to the fundamental limit reminds us that it is possible to alleviate the ADC resolution bottleneck in communication receivers by exploring new quantization mechanisms.
@@ VIII Concluding Remarks
The goal of this study is to evaluate the impact of resolution reduction on information-theoretic limits of communications with receiver quantization. Leveraging the GMI as a basic tool, which enables us to take the decoding rule into consideration, we establish an array of exact and asymptotic results under a standard transceiver architecture. Our results indicate a critical issue in system design that arises as the resolution decreases, namely, optimizing the loading factor by gain control, which minimizes the loss in achievable rate by eliminating the loading loss. The remaining irreducible loss is an appropriate evaluation of the inherent robustness of the considered transceiver architecture. Our results also establish explicit connections between the MSE (affected by the gain control) and the rate loss. Although for general receiver quantization rules, smaller MSE does not necessarily imply smaller rate loss, for the commonly used uniform quantizer we prove that the unique loading factor that minimizes the MSE also maximizes the GMI (i.e., achieves zero loading loss). For perfect gain control, we show that the irreducible loss is determined by the product of the normalized MMSE and the SNR, and provide a geometrical interpretation for this result. Performance approximations and per-bit rules in this case are also given. For imperfect gain control, to understand its impact, we characterize the decay of small rate loss in the high-resolution regime, and propose approximations of the achievable rate as a function of the loading factor which is fairly accurate for moderate resolutions. These results provide insight into transceiver design with nonnegligible quantization effect, especially choice of quantizer resolution and design of AGC.
A limitation of this work is that the obtained analytical results are derived under the Gaussian codebook. Like many classical information-theoretic results given by mutual information, for the GMI, it is also not easy to get closed-form or analytical results without assuming Gaussian input. For practical systems with finite alphabet, we can expect that our result applies well for Gaussian-like constellations that comes from constellation shaping. But it is natural to ask whether our results are still useful in the cases of QAM, phase-shift keying (PSK), and other commonly used inputs, which are all bounded (unlike the unbounded Gaussian input). Since in these cases Proposition 1 does not apply, the GMI evaluation may rely heavily on numerical computation. We left this problem to future study. Nevertheless, the insight gained in this work will be very helpful for the finite-alphabet case.
@@ Appendix B Proof of Proposition 5
$\mathsf{mse}=\;$ $1\displaystyle 2\sum\limits_{k=1}^{K}\int_{\ell_{k-1}}^{\ell_{k}}y_{k}^{2}\phi(v)\mathrm{d}v-2\sum\limits_{k=1}^{K}\int_{\ell_{k-1}}^{\ell_{k}}2y_{k}v\phi(v)\mathrm{d}v+$ 1 (160a)
=\;= $1-4\sum\limits_{k=1}^{K}y_{k}\left(\phi(\ell_{k-1})-\phi(\ell_{k})\right)+2\sum\limits_{k=1}^{K}y_{k}^{2}\left(Q(\ell_{k-1})-Q(\ell_{k})\right)$ (160b)
=\;= $1-\frac{2}{\pi}\left(\sqrt{2\pi}\mathcal{A}-\mathcal{B}\right).$ (160c)
@@
Appendix C Proof of Proposition 6
According to (21), we can derive (44) by straightforward derivation as
$\mathcal{A}=\;$ $\sum\limits_{k=1}^{K-1}\left(k\ell-\frac{\ell}{2}\right)\left(\exp\frac{-(k-1)^{2}\ell^{2}}{2}-\exp\frac{-k^{2}\ell^{2}}{2}\right)+\left(K\ell-\frac{\ell}{2}\right)\exp\frac{-(K-1)^{2}\ell^{2}}{2}$ (162a)
=\;= $\sum\limits_{k=1}^{K-1}k\ell\left(\exp\frac{-(k-1)^{2}\ell^{2}}{2}-\exp\frac{-k^{2}\ell^{2}}{2}\right)-\frac{1}{2}\sum\limits_{k=1}^{K-1}\ell\left(\exp\frac{-(k-1)^{2}\ell^{2}}{2}-\exp\frac{-k^{2}\ell^{2}}{2}\right)$
@@ Appendix C Proof of Proposition 6
=\;= $\sum\limits_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}-\frac{\ell}{2},$ (162c)
@@
=\;= $\pi\ell^{2}\bigg{(}\sum\limits_{k=1}^{K-1}k^{2}\left(Q((k-1)\ell)-Q(k\ell)\right)-\sum\limits_{k=1}^{K-1}k\left(Q((k-1)\ell)-Q(k\ell)\right)+\sum\limits_{k=1}^{K-1}\frac{1}{4}\left(Q((k-1)\ell)-Q(k\ell)\right)\bigg{)}$
$+\pi\left(K-\frac{1}{2}\right)^{2}\ell^{2}Q((K-1)\ell)$ (163b)
=\;= $\pi\ell^{2}\left(\sum\limits_{k=0}^{K-1}(2k+1)Q(k\ell)-\sum\limits_{k=0}^{K-1}Q(k\ell)+\frac{1}{8}\right)$ (163c)
=\;= $\pi\sum\limits_{k=0}^{K-1}2k\ell^{2}Q(k\ell)+\frac{1}{8}\pi\ell^{2}.$ (163d)

### v2

@@ VII-B On Improving Uniform Quantization by Post-Processing
In [58, 40], it has been shown that the GMI (14) in Proposition 1 can be improved by introducing a post-processing of the channel output as $Y\to\tilde{Y}$, thereby extending the channel $X\to Y$ to $X\to\tilde{Y}$; The GMI is maximized by letting
@@
i.e., using a conditional expectation operator as the post-processor, which gives an MMSE estimation of $X$ under the observation $Y$. Utilizing this result, we may improve a uniform quantizer by changing its levels. In fact, the changed levels, denoted by $\{y^{\textrm{c}}_{k}\}$, can be obtained analytically by its step size $\ell$ as [31]
$y^{\textrm{c}}_{k}=\frac{\phi((k-1)\ell)-\phi(k\ell)}{Q((k-1)\ell)-Q(k\ell)},\;k=1,...,K-1$ (150)
@@
$y^{\textrm{c}}_{K}=\frac{\phi((K-1)\ell)}{Q((K-1)\ell)}.$ (151)
Such levels satisfy $y^{\textrm{c}}_{k}=\mathrm{E}\left[\frac{V}{\sigma_{v}}|\frac{V}{\sigma_{v}}\in[(k-1)\ell,k\ell)\right],\;k=1,...,K-1$ and $y^{\textrm{c}}_{K}=\mathrm{E}\left[\frac{V}{\sigma_{v}}|\frac{V}{\sigma_{v}}\geq(K-1)\ell\right]$, where $V\sim\mathcal{N}(0,\sigma_{v}^{2})$ is the input of the quantizer; i.e., the optimal level is the centroid of the corresponding interval. This is consistent with the aforementioned result that a conditional expectation operator maximizes the GMI. In particular, the optimized quantizer with equispaced thresholds considered in Sec. VII-A (see Fig. 17(e)) is equivalent to a uniform quantizer with the same step size $\ell^{\circ}=0.5646$ (found by numerical optimization) combined with an MMSE post-processor which changes mid-rise levels to centroids. Comparing Fig. 17(e) and Fig. 17(d), we see that such a post-processing indeed improves the maximum GMI slightly. Interestingly, the optimal step size $\ell^{\circ}$ that maximizes the GMI in Fig. 17(e) is smaller than the optimal step size $\ell^{*}=0.5860$ in Fig. 17(d) (see also Table I). Thus, when post-processing is available, the step size that maximizes the GMI is reduced. Consequently, a concatenation of an $\ell^{*}$-spaced uniform quantizer and an MMSE post-processor is strictly worse than a single $\ell^{\circ}$-spaced optimized quantizer.
Since the example Fig. 17(e) shows that the gain of the MMSE post-processing is limited for low-resolution uniform quantization, we may infer that the gain is also limited under high resolution. In fact, as the step size decreases, the centroid of the interval $(k-1)\ell,k\ell)$ given in (150) tends to be its midpoint. [^12] In [55] the asymptotic convergence of centroids to midpoints was proved under uniform quantization with infinitely many levels. To see this, note that for an arbitrary $c>\ell/2$, we have
@@
$=\;$ $\lim\limits_{\ell\to 0}\frac{\frac{c-\frac{\ell}{2}}{2}\phi\left(c-\frac{\ell}{2}\right)+\frac{c+\frac{\ell}{2}}{2}\phi\left(c+\frac{\ell}{2}\right)}{\frac{1}{2}\left(\phi\left(c+\frac{\ell}{2}\right)+\phi\left(c-\frac{\ell}{2}\right)\right)}$ (152b)
$=\;$ $c,$ (152c)
which follows from L’Hôpital’s rule. Moreover, according to (101), the largest centroid $y^{\textrm{c}}_{K}$ given in (151) satisfies
$(K-1)\ell>y^{\textrm{c}}_{K}>(K-1)\ell+\frac{1}{(K-1)\ell},$ (153)
implying that $y^{\textrm{c}}_{K}/\ell$ converges to K1K-1 as $\ell\to 0$. From these facts we can conclude that the gain of MMSE post-processing asymptotically vanishes as the resolution increases.
VII-C On Possible Gain of Further Quantization Rules
@@
As shown in Fig. 17(b) and Fig. 17(e), for a nonuniform quantizer, if all the thresholds and levels, rather than only a single factor ($g$ or $L$), are optimized (possibly under a constraint that the thresholds are equispaced), then there is also a consistency between GMI maximization and MSE minimization. In [References, Appendix E] a proof of the consistency has been given (the proof goes through under a equispaced-thresholds constraint). Thus, the maximum achievable rate can be expressed as
@@
where $\mathsf{mmse}^{\textrm{nu}}$ is the normalized minimum MSE of all possible nonuniform quantizers. This result can also be understood via a similar geometrical interpretation as that given in Sec. IV-C. From (154) we can infer that, under our setting of transceiver architecture, the maximum gain in achievable rate from replacing the optimal uniform quantizer by the optimal nonuniform quantizer is determined by the reduction of the normalized MMSE from $\mathsf{mmse}$ to $\mathsf{mmse}^{\textrm{nu}}$. The numerical results in Fig. 17(a) and Fig. 17(b) show that the rate gain is marginal for $b=2$. For higher resolutions the rate gain grows slowly. When $b=4$, according to the numerical results in [31], the optimal non-uniform quantization improves the saturation rate $\log(1/\gamma)$ by no more than $4.4\%$, and optimized quantization with equispaced thresholds improves the saturation rate by $1.5\%$. In fact, for nonuniform quantization with Gaussian input, the asymptotic quantization theory, in particular the Panter-Dite formula [74], yields a “6-dB-per-bit” rule [^13] There is another version of the $6$-dB-per-bit rule as $B_{\mathrm{eff}}=(\mathsf{SNR}_{q}-1.76)/6.02$, which has been widely used in practice to calculate the effective number of bits (ENoB) of a quantizer [2, 3]. It is derived under assumptions that the granular distortion is uniformly distributed over $\ell/2,\ell/2]$ and there is no overload distortion; see [2] for more details. for the maximum SNR as [42]
@@
where $\mathsf{SNR}_{q}^{\textrm{nu}}=1/\mathsf{mmse}^{\textrm{nu}}$. Comparing with (141), we can infer that under our setting the rate gain is limited in the range of resolution of interest. In particular, the gain in maximum saturation rate (65) is $\log\left(\mathsf{SNR}_{q}^{\textrm{nu}}/\mathsf{SNR}_{q}\right)$, which increases logarithmically with $b$.
According to (155), even if the best nonuniform scalar quantization is used, the MSE achieved in the high-resolution regime is still $2.72$ times larger than the theoretical limit implied by the rate-distortion function of a Gaussian source [43]. [^14] The comparison is made under the assumption that the signal at the receiver front-end is represented by $b$ bits per channel use before feeding into the decoder. The MSE of the representation can be reduced by lossy source coding (treating the received signal as a source). A scalar quantizer is a very simple lossy source coding scheme. If we allow coded uniform scalar quantization (i.e., representing the quantizer output by a variable-rate lossless code), the gap can be reduced to only $1.53$ dB ($\pi e/6$) worse than the theoretical limit (see [42] and references therein). Unfortunately, coded quantization does not help in communication receivers since the bottleneck therein is the limited resolution of the quantizer rather than the cost of representing its output. Vector quantization may further reduce the gap, and it is possible to extend our analysis to vector quantizers. [^15] Note that the two quantizers we considered in the basic model (4) are equivalent to a 2-dimensional vector quantizer with a rectangular grid of quantization regions. Replacing it by a polar quantizer [42] may achieve better performance. But applying vector quantization to communication receivers is a more challenging topic. Nevertheless, the gap to the fundamental limit reminds us that it is possible to alleviate the ADC resolution bottleneck in communication receivers by exploring new quantization mechanisms.
A final remark is about dithering, an intentional randomization technique in source quantization [42], which may also be beneficial in receiver quantization; see, e.g., [32], which showed that nonsubtractive Gaussian dithering [^16] In nonsubtractive dithering, one adds a dither signal $W_{\mathrm{d}}$ independent of the quantizer input $V$ before quantization, yielding an output $q(V+W_{\mathrm{d}})$. Subtractive dithering has an additional step that subtracts the dither signal after quantization and finally obtains $q(V+W_{\mathrm{d}})-W_{\mathrm{d}}$. See [42] for more discussion. improves GMI under certain circumstances. In our setting, although nonsubstractive dithering reduces the SNR, it also changes $\gamma$ by enlarging the standard deviation of the input so that all normalized thresholds and levels are scaled by a factor smaller than one. The overall impact on GMI is thus unclear because it depends on the change of the product $\gamma\cdot\mathsf{SNR}$. Nevertheless, for uniform quantization, if we employ perfect gain control which guarantees $\gamma\equiv\mathsf{mmse}$, then nonsubstractive Gaussian dithering is always harmful since the GMI (54) satisfies
$\frac{\mathrm{d}I^{*}_{\textrm{GMI}}}{\mathrm{d}\mathsf{SNR}}=\frac{1-\mathsf{mmse}}{(1+\mathsf{SNR})(1+\mathsf{mmse}\mathsf{SNR})}>0.$ (156)
@@ VIII Concluding Remarks
The goal of this study is to evaluate the impact of resolution reduction on information-theoretic limits of communications with receiver quantization. Leveraging the GMI as a basic tool, which enables us to take the decoding rule into consideration, we establish an array of exact and asymptotic results under a standard transceiver architecture. Our results indicate a critical issue in system design that arises as the resolution decreases, namely, optimizing the loading factor by gain control, which minimizes the loss in achievable rate by eliminating the loading loss. The remaining irreducible loss is an appropriate evaluation of the inherent robustness of the considered transceiver architecture. Our results also establish explicit connections between the MSE (affected by the gain control) and the rate loss. Although for general receiver quantization rules, smaller MSE does not necessarily imply smaller rate loss, for the commonly used uniform quantizer we prove that the unique loading factor that minimizes the MSE also maximizes the GMI (i.e., achieves zero loading loss). For perfect gain control, we show that the irreducible loss is determined by the product of the normalized MMSE and the SNR, and provide a geometrical interpretation for this result. Performance approximations and per-bit rules in this case are also given. For imperfect gain control, to understand its impact, we characterize the decay of small rate loss in the high-resolution regime, and propose approximations of the achievable rate as a function of the loading factor which are fairly accurate for moderate resolutions. These results provide insight into transceiver design with nonnegligible quantization effect, especially regarding the choice of quantizer resolution and the design of AGC.
A limitation of this work is that the obtained analytical results are derived under the Gaussian codebook. Like many classical information-theoretic results given by mutual information, for the GMI, it is also not easy to get closed-form or analytical results without assuming Gaussian input. For practical systems with finite alphabet, we can expect that our result applies well for Gaussian-like constellations that come from constellation shaping. But it is natural to ask whether our results are still useful in the cases of QAM, phase-shift keying (PSK), and other commonly used inputs, which are all bounded (unlike the unbounded Gaussian input). Since in these cases Proposition 1 does not apply, the GMI evaluation may rely heavily on numerical computation. We left this problem to future study. Nevertheless, the insight gained in this work will be very helpful for the finite-alphabet case.
@@ Appendix B Proof of Proposition 5
$\mathsf{mse}=\;$ $2\sum\limits_{k=1}^{K}\int_{\ell_{k-1}}^{\ell_{k}}y_{k}^{2}\phi(v)\mathrm{d}v-2\sum\limits_{k=1}^{K}\int_{\ell_{k-1}}^{\ell_{k}}2y_{k}v\phi(v)\mathrm{d}v+1$ (160a)
$=\;$ $1-4\sum\limits_{k=1}^{K}y_{k}\left(\phi(\ell_{k-1})-\phi(\ell_{k})\right)+2\sum\limits_{k=1}^{K}y_{k}^{2}\left(Q(\ell_{k-1})-Q(\ell_{k})\right)$ (160b)
$=\;$ $1-\frac{2}{\pi}\left(\sqrt{2\pi}\mathcal{A}-\mathcal{B}\right).$ (160c)
@@
Appendix C Proof of Corollary 6
According to (23), we can derive (43) straightforwardly as
[…]
$=\;$ $\sum\limits_{k=0}^{K-2}(k+1)\ell\exp\frac{-k^{2}\ell^{2}}{2}-\sum\limits_{k=1}^{K-1}k\ell\exp\frac{-k^{2}\ell^{2}}{2}-\sum\limits_{k=0}^{K-2}\frac{\ell}{2}\exp\frac{-k^{2}\ell^{2}}{2}+\sum\limits_{k=1}^{K-1}\frac{\ell}{2}\exp\frac{-k^{2}\ell^{2}}{2}$
@@ Appendix C Proof of Corollary 6
$=\;$ $\sum\limits_{k=0}^{K-1}\ell\cdot\exp\frac{-k^{2}\ell^{2}}{2}-\frac{\ell}{2},$ (162c)
@@
$=\;$ $\pi\ell^{2}\Bigg(\sum\limits_{k=1}^{K-1}k^{2}\left(Q((k-1)\ell)-Q(k\ell)\right)-\sum\limits_{k=1}^{K-1}k\left(Q((k-1)\ell)-Q(k\ell)\right)+\sum\limits_{k=1}^{K-1}\frac{1}{4}\left(Q((k-1)\ell)-Q(k\ell)\right)$
$+\left(K^{2}-K+\frac{1}{4}\right)Q((K-1)\ell)\Bigg)$ (163b)
$=\;$ $\pi\ell^{2}\Bigg(\sum\limits_{k=0}^{K-2}(k+1)^{2}Q(k\ell)-\sum\limits_{k=0}^{K-1}k^{2}Q(k\ell)-\sum\limits_{k=0}^{K-2}(k+1)Q(k\ell)+\sum\limits_{k=0}^{K-1}kQ(k\ell)+\frac{1}{4}\sum\limits_{k=0}^{K-2}Q(k\ell)-\frac{1}{4}\sum\limits_{k=0}^{K-1}Q(k\ell)+\frac{Q(0)}{4}$
$+\left(K^{2}-K+\frac{1}{4}\right)Q((K-1)\ell)\Bigg)$ (163c)
$=\;$ $\pi\ell^{2}\Bigg(\sum\limits_{k=0}^{K-1}(2k+1)Q(k\ell)-K^{2}Q((K-1)\ell)-\sum\limits_{k=0}^{K-1}Q(k\ell)+KQ((K-1)\ell)-\frac{1}{4}Q((K-1)\ell)+\frac{1}{8}$
$+\left(K^{2}-K+\frac{1}{4}\right)Q((K-1)\ell)\Bigg)$ (163d)
$=\;$ $\pi\sum\limits_{k=0}^{K-1}2k\ell^{2}Q(k\ell)+\frac{1}{8}\pi\ell^{2}.$ (163e)